OpenStudy (anonymous):

Find the limit of the function algebraically. lim(x -> 5) x^2-25/x-5

myininaya (myininaya):

So I think you mean (x^2-25)/(x-5) right? This expression does not exist at x=5 because when you plug in 5 you get something/0 Now when you have 0/0 that means the limit can exist Is there any algebra you can perform to force a function that is continuous at 5?

myininaya (myininaya):

Basically is there any factoring you can do?

OpenStudy (anonymous):

Yes that's it and simplify iit

myininaya (myininaya):

yep then plug in 5 to resulting function that is continuous at 5

myininaya (myininaya):

you can share your steps with me if you want

OpenStudy (anonymous):

would it be 10/0?

myininaya (myininaya):

not 10/0 nope show me what you did to get that

OpenStudy (anonymous):

(x+25)(x-25)/x-5

myininaya (myininaya):

that isn't how you factor x^2-25

myininaya (myininaya):

x^2-a^2=(x-a)(x+a) you have x^2-5^2=(x-5)(x+5)

OpenStudy (anonymous):

OH (x+5)(x-5)

myininaya (myininaya):

right but what do yo notice about the top and bottom (As far as factors go)

myininaya (myininaya):

and you have (x+5)(x-5)/(x-5)

OpenStudy (anonymous):

You can cancel

OpenStudy (anonymous):

(x+5)

myininaya (myininaya):

right you basically get rid of the discontinuity so you can plug in x=5

OpenStudy (anonymous):

so 10

myininaya (myininaya):

right

OpenStudy (anonymous):

Thank you!!

myininaya (myininaya):

np